# Capacity-Aware Inference: Mitigating the Straggler Effect in Mixture of Experts

### v3

@@ Appendix A Implementation Details
To evaluate model performance, we report normalized zero-shot or few-shot accuracy on the LM-Harness benchmark. The number of shots for each task is detailed in Table 4, which includes multiple tasks: ARC-C [7], BoolQ [6], HellaSwag [43], MMLU [20], OBQA [30], PIQA [2], RTE [40], WinoGrande [1] and GSM8K [8]. The evaluation code is based on EleutherAI’s LM Harness framework [16].
@@
Appendix B Multi-Modal Tasks
@@
As in MME and MMBench, the Image-First variants of Token Drop and Expanded Drop also exhibit consistent effectiveness on SEED-Bench [27], maintaining strong performance even under low capacity factors such as . In addition to the redundancy in image tokens, Figure 10 shows that text tokens constitute only a small portion of the total token assignments. This allows the regulation of to retain almost all text tokens under the Image-First strategy.
@@
Appendix C Pseudocode for Token Drop and Expanded Drop
We present the detailed pseudo-code for the algorithm implementations, as shown in Algorithm 1 and Algorithm 2. Both of these two methods adopt the selection metrics of “Score.”
Appendix D Maximum Expert Selection
Expanded Drop employs overselection and dropping to not only ensure load balance but also improve the utilization of underloaded experts. Although this process may allow some tokens to select more than experts, Table 7 shows that enforcing a strict -expert limit is unnecessary: allowing additional expert selections has the potential of enhancing representational capacity, while rigid constraints introduce redundant computations.
@@ Appendix D Maximum Expert Selection
Appendix E Layer-wise Expert Load
To analyze imbalanced token assignments, we measure the expert load for each expert by tracking the peak expert load while running MoE models on various test datasets. Figure 12, 13, 14 and 15 present the full results for the normalized layer-wise expert load for OLMoE, DeepSeek-V2, Qwen1.5-MoE, and Mixtral-87B-Instruct, respectively.
Appendix F Calculation Dropped tokens
Based on Equation 12, we calculate the total number of dropped tokens across experts in each layer under different capacity factors, as illustrated in Figures 16, 18, 17, and 19.

### v4

@@ Appendix A Implementation Details
We mainly focus on lightweight MoE models (less than 20B parameter budget). We conduct experiments on OLMoE (Muennighoff et al., 2024), Qwen1.5-MoE (Team, 2024b), DeepSeek-V2-Lite (et al., 2024a), Mixtral (Jiang et al., 2024), MolmoE (Deitke et al., 2024), and Qwen3-MoE (Yang et al., 2025), due to their competitive performance and widespread applications.
@@
To evaluate model performance, we report normalized zero-shot or few-shot accuracy on the LM-Harness benchmark. The number of shots for each task is detailed in Table 5, which includes multiple tasks: ARC-C (Clark et al., 2018), BoolQ (Clark et al., 2019), HellaSwag (Zellers et al., 2019), MMLU (Hendrycks et al., 2021), OBQA (Mihaylov et al., 2018a), PIQA (Bisk et al., 2019), RTE (Wang et al., 2019), WinoGrande (44), and GSM8K (Cobbe et al., 2021). The evaluation code is based on EleutherAI’s LM Harness framework (Gao et al., 2023).
Appendix B Additional Results on Language Tasks
Long-context understanding poses substantial challenges for language models. We therefore investigate whether Token Drop and Expanded Drop can maintain competitive performance in this setting. Table 6 presents their results on a long in-context learning benchmark. Following LongICLBench (Li et al., 2024b), we evaluate on the BANKING77 dataset (Casanueva et al., 2020), with the only modification being that we sample 100 test examples. BANKING77 is an intent detection dataset in the banking domain with 77 classes. We consider 1-shot/label to 5-shot/label settings, corresponding to context lengths of 2K, 4K, 7K, 9K, and 14K tokens. Across different input sequence lengths, both Token Drop and Expanded Drop consistently match the performance of the baseline. This demonstrates the strong generalization ability and wide applicability of these techniques.
| Method | 2K | 4K | 7K | 9K | 11K | Avg. |
| Baseline | 0.28 | 0.44 | 0.56 | 0.62 | 0.70 | 0.52 |
| Token Drop | 0.27 | 0.43 | 0.57 | 0.65 | 0.71 | 0.53 |
| Expanded Drop | 0.23 | 0.53 | 0.63 | 0.62 | 0.70 | 0.54 |
To further access the robustness of Token Drop and Expanded Drop in generation tasks, we test additional benchmarks such as HumanEval (Chen et al., 2021), NQ-Open (Kwiatkowski et al., 2019), and MTS-Dialog (Ben Abacha et al., 2023) in Table 7.
| Method | OLMoE-Instruct | Qwen1.5-MoE-Chat | DeepSeek-V2-Lite-Chat | ||||||
| HE | NQ | MTS | HE | NQ | MTS | HE | NQ | MTS | |
| Baseline | 0.311 | 0.175 | 0.827 | 0.280 | 0.226 | 0.874 | 0.506 | 0.265 | 0.830 |
| Token Drop | 0.299 | 0.159 | 0.823 | 0.281 | 0.214 | 0.869 | 0.494 | 0.252 | 0.831 |
| Expanded Drop | 0.303 | 0.170 | 0.829 | 0.291 | 0.224 | 0.872 | 0.508 | 0.268 | 0.831 |
The consistent performance on code generation, open-domain question answering, and dialog creation demonstrates the robustness of Token Drop and Expanded Drop for long-range generation tasks.
Appendix C Additional Results on Multi-Modal Tasks
@@
As in MME and MMBench, the Image-First variants of Token Drop and Expanded Drop also exhibit consistent effectiveness on SEED-Bench (Li et al., 2024a), maintaining strong performance even under low capacity factors such as . In addition to the redundancy in image tokens, Figure 11 shows that text tokens constitute only a small portion of the total token assignments. This allows the regulation of to retain almost all text tokens under the Image-First strategy.
@@
Appendix D Ablation Study
Model-Specific Imbalanced Property
We explore the imbalance property in various models, such as OLMoE, DeepSeek, Qwen and Mixtral, which differ in both architectures (e.g., depth and width) and training strategies (e.g., training from scratch (Muennighoff et al., 2024; et al., 2024a) vs. training after upcycling (Jiang et al., 2024; Team, 2024b)).
On the one hand, our findings in Appendix F reveal different training strategies result in significantly varying levels of imbalance. Specifically, MoE models trained from scratch exhibit a much higher degree of imbalance. For instance, OLMoE and DeepSeek-V2-Lite experience peak expert-wise token allocations exceeding , whereas Qwen1.5-MoE and Mixtral are upcycled from dense language models, maintain a more balanced distribution, with peak expert-wise allocations staying below . This is because upcycling initializes all experts with identical parameters (Komatsuzaki et al., 2023), reducing divergence and promoting balanced training in the early stages.
Capacity Factor
Speedup Measurement Under Different Workloads
To assess the effectiveness of Token Drop in efficiency, we measure speedup under a range of token workloads. Specifically, we vary both the sequence length and batch size, and present the speedup results on DeepSeek-V2 with a capacity factor of 2.0 in Table 10.
| Batch Size | 8K | 8K | 8K | 4K | 2K | 2K | 1K | 1K | 1K |
| Prompt Length | 0.1K | 0.2K | 0.4K | 1K | 1K | 2K | 1K | 2K | 4K |
| Speedup |
The straggler effect becomes more pronounced under heavier workloads, where GPUs operate at higher utilization with limited spare capacity, making the speedup more noticeable. In practical server-side MoE deployments, workloads are substantially higher, and our techniques effectively mitigate the resulting straggler effect.
Maximum Number of Selected Experts
Expanded Drop adopts an overselection-and-dropping strategy that not only maintains load balance but also improves the utilization of underloaded experts. Although this mechanism allows some tokens to select more than experts, Table 11 demonstrates that such flexibility benefits downstream task performance, suggesting that enforcing a strict maximum of experts is unnecessary. Allowing additional expert selections can enhance representational capacity, whereas rigid constraints may unnecessarily limit model performance.
@@ Appendix D Ablation Study
Appendix E Implementation-level Code for Token Drop and Expanded Drop
The detailed implementation-level code for Token Drop and Expanded Drop is presented in Algorithms 1 and 2, respectively.
Appendix F Layer-wise Expert Load
To analyze imbalanced token assignments, we measure the expert load for each expert by tracking the peak expert load while running MoE models on various test datasets. Figure 13, 14, 15 and 16 present the full results for the normalized layer-wise expert load for OLMoE, DeepSeek-V2, Qwen1.5-MoE, and Mixtral-87B-Instruct, respectively.
Appendix G Calculation Dropped tokens
Based on Equation 11, we calculate the total number of dropped tokens across experts in each layer under different capacity factors, as illustrated in Figures 17, 19, 18, and 20.
![[Uncaptioned image]](figs/example.jpg)